# Locomotion on Constrained Footholds via Layered Architectures and Model Predictive Control

### v1

@@ I-A Related Works
Figure 2: Diagram of the control architecture. Velocity commands from the user are used for sampling and go into the MPC cost. A set of contacts (terrain surfaces) are sampled and passed into the MPC as polytope constraints. The MPC computes full order trajectories in parallel for each set of polytopes, selecting the best to use. The MPC then outputs feedforward torques, $\tau$, joint configurations, $q$, and joint velocities $v$ as shown in the middle box. In the MPC box, the feet trajectories are shown in a green-red gradient where green is closer in time. The polytopes to be stepped on are shown highlighted in yellow, white, and pink (color coded to each foot). The blue lines show the force computed by the MPC. A state estimator outputs the state at 500 Hz.
@@
Contact Implicit MPC (or CIMPC), attempts to solve for contacts and motion in a single, gradient based optimization. There have been a number of approaches to this problem using a variety of approximations and optimization techniques. Some works have used smooth, non-physical, ground reaction forces [11, 12]. On the other hand, a number of methods have used hard contacts [13, 14, 15, 16]. Alternatively, a gradient based method with optimization based contact models was proposed in [17]. In general, these solvers are subject to poor numerical conditioning [18] and slower solve speeds than fixed contact MPC.
@@ I-B Contributions
In this work, we propose a layered architecture with MPC to locomote through complex terrain. The layered architecture has two components: (1) a sampling-based, gradient-free method for selecting terrain for locomotion and, (2) a full order nonlinear MPC controller outputting PD setpoints and feed forward torques. The layered control architecture allows the robot to select which footholds to use in real time while considering the closed loop performance of the system for a variety of terrain options.
The controller is faster than pure sampling-based and CIMPC methods, and it can be run in real time for a humanoid. The proposed controller also outperforms the Raibert heuristic [24] for selecting terrain when traveling at high speeds, as shown in simulation experiments. Further, the controller does not require heuristics, which allows it to generalize to other tasks beyond stepping stones.
We demonstrate the performance of the controller on the Unitree Go2 quadruped robot crossing over a variety of terrains and locomoting on flat ground, as shown in Fig. 1. Simulation results shown the efficacy of the approach on the Unitree G1 humanoid crossing over complex terrain.
@@ II Layered Control and MPC Formulation
The architecture has two major parts: a gradient free algorithm that chooses the discrete variables (i.e. selects which pieces of terrain to use), and a smooth MPC that fixes these discrete variables and is formulated as a classical smooth Nonlinear Program (NLP). We choose the discrete varaibles through a sampling based method: given the velocity command to the robot, we propagate forward a nominal stance foot location and sample the terrain around that location. Then, the fixed mode MPC is solved in parallel for every sample. The lowest cost MPC solve is used as the controller output. To enforce the terrain sample in the fixed mode MPC, we use a polytope constraint on the end effector during stance, similar to [10]. The entire architecture is shown in Fig. 2.
@@ II-A1 Dynamics Constraints
$\delta v_{k+1}\approx-\nabla_{\delta v_{k+1}}\bar{f}^{-1}(\nabla_{\delta q_{k}}\bar{f}\delta q_{k}+\nabla_{\delta v_{k}}\bar{f}\delta v_{k}+\\ \nabla_{\delta F_{k}}\bar{f}\delta F_{k}-\bar{\tau}_{k}-\delta\tau_{k}+\bar{f}).$
@@ II-A2 Inequality Constraints
$\gamma_{k}^{l}A^{l}_{k}r^{l}(\bar{q}_{k}\oplus\delta q_{k})\leq\gamma_{k}^{l}b^{l}_{k}.$
where we use the notation $r^{l}(q)$ to denote the global 3-D position (determined via forward kinematics) of end effector $l$. $A$ and $b$ can be different at each node for each end effector and thus have the corresponding sub and superscripts.
@@ II-A3 Equality Constraints
$\gamma_{k}^{l}r^{l}(\bar{q}_{k}\oplus\delta q_{k},\bar{v}_{k}+\delta v_{k})=0.$
@@ II-A4 Cost Function
$H_{x}$ $\approx\nabla_{\delta q_{k}}(r^{l}(\bar{q}_{k}\oplus\delta q_{k}))^{T}W_{r}\nabla_{\delta q_{k}}(r^{l}(\bar{q}_{k}\oplus\delta q_{k})).$
@@ II-B Sampling for Gradient Free Decision Making
To select terrain for locomotion, we use a sampling based method where we evaluate the cost of stepping on each piece of terrain and choose the “best” one. To do this, we parallelize the MPC computation and choose different polytopes for each MPC solve. Note that this same idea can be applied to the optimizing the contact schedule where $\gamma$ is decided at each node, but in this work we focus on the sampling of terrain. The specifics of the sampling algorithm presented below are thus tailored to that application.
The sampling algorithm works as follows. First, given a desired velocity, we propagate the torso of the robot forward in the $x$-$y$ plane. Given the resulting torso placement, we check a nominal foot location, i.e., right below the hip at the middle of the stance phase [24]. If the nominal foot location is not over a piece of terrain, then the sampling process continues; otherwise, we use the terrain at the nominal location [^1] Technically, you could choose to sample here even if it is over a piece of terrain (and thus not use the nominal position heuristic), but to reduce the sample space, we choose to take the nominal terrain..
When multiple legs have the same contact phases, then we choose to consider their sampling together to prevent re-sampling the same terrain combinations. A “sample tree” is used to keep track of the possibilities. This tree holds all possible combinations of samples for a single contact phase. The nodes of the tree represent discrete pieces of terrain. To understand the tree structure better, consider an example where there are two legs with the same contact phases. In this case, the tree will have a depth of three: a dummy root node with a number of children equal to the number of terrain options for the first foot, then each of those nodes has a number of children for all the terrain options for the second foot. Therefore, a given branch of the tree corresponds to a valid, unique, terrain selection.
To create the tree and prevent sampling terrain that is outside of the kinematic limits of the robot, we create a circle in the $x$-$y$ plane centered on the nominal foot location that approximates the kinematic limits of the robot [^2] Note that any method that gives a true or approximate representation of the kinematic limits can be used here. We find it easy and effective to use a circle about the nominal position.. Then we calculate the area of the intersection of each piece of terrain within the circle. Any piece of terrain with non-zero area intersection is added as a node.
@@
number of MPC samples $M$, number of contact phases for each MPC $L$, nominal feet positions to be sampled around $\eta_{i}\in\mathbb{R}^{2K}\quad\forall i\leq L$ with $K$ the number of feet in that phase
@@
         $\triangleright$ More samples than unique terrain options
@@
Figure 3: Example result from the sampling. At sampling time, all feet are on the left-most tile (signified by the large circles), and we are plotting all 4 sampled paths. The plotted lines show which blocks are sampled, and the plotted position of the circles do not correspond with the feet location within the block.
@@ IV Results
Fig. 5 shows a hardware experiment with the Unitree Go2 quadruped. The quadruped needs to traverse a gap where there are a few small cinder blocks to step on. Here we use four samples and see that the quadruped can successfully traverse the terrain. This demonstrates that the policy can run in the real world and is robust to model-mismatches. Beyond trivial model mismatch, we noticed that the large wooden platforms were not completely stationary and would wobble, thus providing other real-world disturbances. In addition, the lidar unit mounted on the quadruped was not modeled in the MPC (the lidar weight is about 265 g). Fig. 1 shows snapshots from a few other hardware experiments.
@@
Robot Optimization DoF Nodes # Parallel MPC Prep Time (s) Feedback Time (s)
@@ V Conclusion And Future Work
The examples in this paper were specifically focused on the stepping stone problem, but we believe the presented architecture to be more broad than that. Future work should focus on using this for other tasks, such as using a hand to brace against a wall in push recovery, or optimizing the contact schedule. Further, there is lots of room for experimentation with the sampling algorithm. We chose to sample directly at the terrain layer, but one could imagine that the sampling layer could be done with a full order optimizer like MPPI and the relevant contact information extracted out and passed to the gradient based MPC.
@@
We presented a layered controller with MPC for dynamic locomotion over constrained footholds. This controller demonstrates that through the use of both gradient-free and gradient-based methods, both discrete and continuous variables can be optimized in real time to improve performance. The controller preserves the speed of MPC while getting the optimality and generality of sampling based methods. We demonstrated these capabilities with hardware and simulation experiments.

### v2

@@ I-A Related Works
Figure 2: Diagram of the control architecture. Velocity commands from the user are used for sampling and go into the MPC cost. A set of contacts (terrain surfaces) are sampled and passed into the MPC as polytope constraints. The MPC computes full order trajectories in parallel for each set of polytopes, selecting the lowest cost option to use. The MPC then outputs feedforward torques, $\tau$, joint configurations, $q$, and joint velocities $v$ as shown in the middle box. In the MPC box, the feet trajectories are shown in a green-red gradient where green is closer in time. The polytopes to be stepped on are shown highlighted in yellow, white, and pink (color coded to each foot). The blue lines show the force computed by the MPC. A state estimator outputs the state at 500 Hz.
@@
Contact Implicit MPC (or CIMPC), attempts to solve for contacts and motion in a single, gradient based optimization. There have been a number of approaches to this problem using a variety of approximations and optimization techniques. Some works have used smooth, non-physical, ground reaction forces [11, 12]. On the other hand, a number of methods use hard contacts [13, 14, 15, 16]. Alternatively, a gradient based method with optimization based contact models was proposed in [17]. In general, these solvers are subject to poor numerical conditioning [18] and slower solve speeds than fixed contact MPC.
@@ I-B Contributions
In this work, we propose a layered architecture with MPC to locomote through complex terrain. The layered architecture has two components: (1) a sampling-based, gradient-free method for selecting terrain for locomotion and, (2) a full order nonlinear MPC controller outputting PD setpoints and feed forward torques. The layered control architecture allows the robot to select which footholds to use in real time while considering the closed loop performance of the system for a variety of terrain options. In addition, the MPC is formulated to optimize on the quaternion tangent space rather than using Euler angles as is commonly done in MPC.
The controller is faster than pure sampling-based and CIMPC methods, and it can be run in real time for a humanoid. The proposed controller also outperforms the Raibert heuristic [24] for selecting terrain when traveling at high speeds, as shown in simulation experiments. Further, the controller does not require heuristics, which allows it to generalize to other tasks beyond stepping stones. We demonstrate the performance of the controller on the Unitree Go2 quadruped robot crossing over a variety of terrains and locomoting on flat ground, as shown in Fig. 1. Simulation results shown the efficacy of the approach on the Unitree G1 humanoid crossing over complex terrain.
@@ II Layered Control and MPC Formulation
The architecture has two major parts: a gradient free algorithm that chooses the discrete variables (i.e. selects which pieces of terrain to use), and a smooth MPC that fixes these discrete variables and is formulated as a classical smooth Nonlinear Program (NLP). We choose the discrete variables through a sampling based method: given the velocity command to the robot, we propagate forward a nominal stance foot location and sample the terrain around that location. Then, the fixed mode MPC is solved in parallel for every sample. The lowest cost MPC solve is used as the controller output. To enforce the terrain sample in the fixed mode MPC, we use a polytope constraint on the end effector during stance, similar to [10]. The entire architecture is shown in Fig. 2.
@@ II-A1 Dynamics Constraints
$\delta v_{k+1}\approx-\nabla_{\delta v_{k+1}}\bar{f}^{-1}(\nabla_{\delta q_{k}}\bar{f}\delta q_{k}+\nabla_{\delta v_{k}}\bar{f}\delta v_{k}+\\ \nabla_{\delta F_{k}}\bar{f}\delta F_{k}-\bar{\tau}_{k}-\delta\tau_{k}+\bar{f}).$ (3)
@@ II-A2 Inequality Constraints
$\gamma_{k}^{l}A^{l}_{k}r^{l}(\bar{q}_{k}\oplus\delta q_{k})\leq\gamma_{k}^{l}b^{l}_{k}.$ (5)
where we use the notation $r^{l}(q)$ to denote the global 3-D position (determined via forward kinematics) of end effector $l$. AA and $b$ can be different at each node for each end effector and thus have the corresponding sub and superscripts.
@@ II-A3 Equality Constraints
$\gamma_{k}^{l}r^{l}(\bar{q}_{k}\oplus\delta q_{k},\bar{v}_{k}+\delta v_{k})=0.$ (6)
@@ II-A4 Cost Function
$H_{r}$ $\approx\nabla_{\delta q_{k}}(r^{l}(\bar{q}_{k}\oplus\delta q_{k}))^{T}W_{r}\nabla_{\delta q_{k}}(r^{l}(\bar{q}_{k}\oplus\delta q_{k})).$
@@ II-B Sampling for Gradient Free Decision Making
To select terrain for locomotion, we use a sampling based method where we evaluate the cost of stepping on each piece of terrain throughout the MPC horizon and choose the lowest cost set of terrain. To do this, we parallelize the MPC computation and choose different polytopes for each MPC solve. Note that this same idea can be applied to the optimizing the contact schedule where $\gamma$ is decided at each node, but in this work we focus on the sampling of terrain. The specifics of the sampling algorithm presented below are thus tailored to that application.
The sampling algorithm works as follows. First, given a desired velocity, we propagate the torso of the robot forward in the $x$-$y$ plane throughout the MPC horizon. Given the resulting torso placement, we check a nominal foot location, i.e., right below the hip at the middle of the stance phase [24]. If the nominal foot location is not over a piece of terrain, then the sampling process continues; otherwise, we use the terrain at the nominal location [^1] Technically, you could choose to sample here even if it is over a piece of terrain (and thus not use the nominal position heuristic), but to reduce the sample space, we choose to take the nominal terrain..
When multiple legs have the same contact phases, then we choose to consider their sampling together to prevent re-sampling the same terrain combinations. A “sample tree” is used to keep track of the possibilities. This tree holds all possible combinations of samples for the given set of feet for a single contact phase. The nodes of the tree represent discrete pieces of terrain. To understand the tree structure better, consider an example where there are two legs with the same contact phases. In this case, the tree will have a depth of three (number of feet plus one): a dummy root node with a number of children equal to the number of terrain options for the first foot, then each of those nodes has a number of children for all the terrain options for the second foot. Therefore, a given branch of the tree corresponds to a valid, unique, terrain selection.
To create the tree, and prevent sampling terrain that is outside of the kinematic limits of the robot, we create a circle in the $x$-$y$ plane centered on the nominal foot location that approximates the kinematic limits of the robot [^2] Note that any method that gives a true or approximate representation of the kinematic limits can be used here. We find it easy and effective to use a circle about the nominal position.. Then, we calculate the area of the intersection of each piece of terrain within the circle. Any piece of terrain with non-zero area intersection is added as a node.
@@
number of MPC samples $M$, number of contact phases for each MPC $L$, nominal feet positions to be sampled around, $\zeta_{i}\in\mathbb{R}^{2K}\quad\forall i\leq L$, with $K$ the number of feet in that phase
@@
[…]
@@
Figure 3: Example result from the sampling. At sampling time, all feet are on the left-most tile (signified by the large circles), and we are plotting all 4 sampled paths. The plotted lines show which blocks are sampled. The plotted position of the circles do not correspond with the feet location within the block.
@@ IV Results
Fig. 5 shows a hardware experiment with the Unitree Go2 quadruped. The quadruped needs to traverse a gap where there are a few small cinder blocks to step on. Here we use four samples and see that the quadruped can successfully traverse the terrain. This demonstrates that the controller can run in the real world and is robust to model-mismatches. Beyond trivial model mismatch, we noticed that the large wooden platforms were not completely stationary and would wobble, thus providing other real-world disturbances. In addition, the lidar unit mounted on the quadruped was not modeled in the MPC (the lidar weight is about 265 g). Fig. 1 shows snapshots from a few other hardware experiments.
@@
Robot Optimization DoF Nodes # Parallel MPC Prep Time (ms) Feedback Time (ms)
@@ V Conclusion And Future Work
We presented a layered controller with MPC for dynamic locomotion over constrained footholds. This controller demonstrates that through the use of both gradient-free and gradient-based methods, both discrete and continuous variables can be optimized in real time to improve performance. The controller preserves the speed of MPC while getting the optimality and generality of sampling based methods. We demonstrated these capabilities with hardware and simulation experiments for quadrupeds and bipeds.
The examples in this paper were specifically focused on the stepping stone problem, but we believe the presented architecture to be more broad than that. Future work should focus on using this for other tasks, such as using a hand to brace against a wall in push recovery, or optimizing the contact schedule. Further, there is room for experimentation with the sampling algorithm. We chose to sample directly at the terrain layer, but one could imagine that the sampling layer could be done with a full order optimizer like MPPI and the relevant contact information extracted out and passed to the gradient based MPC.